OpenStudy (anonymous):

Another One , Please Help ? Ill Medal What is the solution to the system of equations? x = 4y + 6 x - 3y = 4 ( -- , -- )

OpenStudy (tkhunny):

Show your work, please. Do NOT post questions with no work shown. Have you considered substitution?

OpenStudy (anonymous):

try plugging (4y+6) for x into x - 3y = 4

OpenStudy (anonymous):

4y+6 for x into x - 3y = 4

OpenStudy (anonymous):

so do 4y - 3y and subtract 6 from both sides =)

OpenStudy (anonymous):

hello ?

OpenStudy (anonymous):

substitute the value 4y+6 and remove the parentheses x=4y+6 4y+6-3y=4 add -3y to 4y and get y , because they are like terms. x=4y+6 y+6=4 you will have to subtract by 6 because 6 doesn't have a variable to solve x=4y+6 y=-6+4 4+-6=-2 for y the value would be -2 for x you will have to x=4(-2)+6 x=-8+6 x=-2 and y =-2

OpenStudy (anonymous):

ill follow rules nxt time =)